In a Math Test, the mean score is 45 and the standard deviation is 4. Assuming normality, what is the probability that a score picked at a random will lie above score 50?


1
Expert's answer
2021-03-01T07:10:08-0500

"P(X>50)=P(Z>\frac{50-45}{4})=P(Z>1.25)=0.1056."